A 1kg mass slides down several frictionless inclines - each incline angle decreases by 1 degree for each 1-meter segment travelled. When the mass reaches the 0-degree incline, determine the total time travel time, the total height the mass falls and velocity.
1
Expert's answer
2020-05-18T10:32:52-0400

If the first meter inclined 90°, the other 89°, and so on, the total height will be:


"H=d\text{ sin}90^\circ+d\text{ sin}89^\circ+...+d\text{ sin}23^\circ+...+d\text{ sin}1^\circ=
\\=d\sum^{90^\circ}_{i=1}\text{sin}n_i=57.79\text{ m}."


Using energy conservation, the final velocity:


"v=\sqrt{2gH}=\sqrt{2\cdot9.8\cdot57.79}=33.7\text{ m/s}."

Ignoring friction and jumps at the edges where inclines are joined, we can try to calculate the time if that mass is falling from height H:


"t=\sqrt{\frac{2H}{g}}=\sqrt{\frac{2\cdot57.79}{9.8}}=3.43\text{ s},"


and compare it with real time calculated according to the expression


"t=\sum^{90}_{i=1}t_i,\\
\space\\
t_i=\frac{v_i+v_{i-1}}{2d}=\frac{v_i+v_{i-1}}{2d},\\
\space\\
v_i=\sqrt{2gd\text{ sin}(90-i)^\circ}.\\
t=33.7\text{ s}."